OpenStudy (ja1):

Which ratio forms a proportion with:

OpenStudy (ja1):

\[\frac{ 27 }{ 105 }\] Simple question, It's just I forgot the rules here xD

OpenStudy (ja1):

Multiple Choice Answers Are: \[\frac{ 9 }{ 37 }\] \[\frac{ 18 }{ 70 }\] \[\frac{ 18 }{ 96 }\] \[\frac{ 18 }{ 63 }\]

OpenStudy (sasogeek):

well find the common factor and cancel out, i would start with 2, if 2 doesn't factor out of both, then i'll move on to 3, until i find a common factor

OpenStudy (ja1):

Ok got it, and so my objective would be to make sure that both the numerator and the denominator are divided by the same number?

OpenStudy (sasogeek):

yes, and once u do that, find out what the result of both the numerator and denominator are when u divide them, the answer will come smiling at you :) trust me

OpenStudy (ja1):

Ah I see, thanks a bunch I had forgotten how to do this xD

OpenStudy (sasogeek):

it's tricky though, don't jump to conclusions after dividing, some multiplication is involved to find the right answer

OpenStudy (ja1):

Ah I see, yah I am usually thorough with my math so I hope I don't mess up lol

OpenStudy (sasogeek):

lemme know what u decide as your answer :)

OpenStudy (ja1):

Hmm....

OpenStudy (ja1):

What I don't get is that they have to be proportionate and 18 does not go into 27...

OpenStudy (sasogeek):

why are we talking about 18 and 27?

OpenStudy (sasogeek):

what did u do?

OpenStudy (ja1):

Well the answers all have 18 as the numerator except for A

OpenStudy (ja1):

I mean the factor I chose was 3, it fot perfectly into both.

OpenStudy (sasogeek):

okay so let me try to help you here :)

OpenStudy (sasogeek):